A child in a boat throws a 5.15-kg package out horizontally with a speed of 10.0 m/s. The mass of the child is 24.2kg and the mass of the boat is 33.2kg .

Calculate the velocity of the boat immediately after, assuming it was initially at rest.
Answer #1

0 = 5.15*10 - (24.2+33.2)v

v = .897 m/sec
Answer #2

mass of entire boat+child = M= 24.2 +33.2

M = 57.4 kg

mass of object m= 5.15 kg

the object velocity is v= 10m/s

initially bothe object,(child+boat)at rest

after throwing object (child+boat) velocity = V (say)

So, here we can say mv + MV =0 (as initially bothe are at rest w.r.t. each other)

57.4 *V + 10* 5.15 = 0

V =- 51.5/ 57.4

V = - 0.8972 m/s

so (child+boat) atmoving velocity V = 0.8972 in oppsite direction of object
Answer #5

using conservation of momentum.
masspackage*velocitypackage+totalmassinboat*velocityboat=0

masspackage*velocitypackage=-totalmassinboat*velocityboat

velocityboat =- (masspackage*velocitypackage)/totalmassinboat


thus ,

velocityboat=-(5.15 * 10.0)/(33.2 + 24.2)

velocityboat= -0.897 m/s

negative sign is for the opposite direction .